Kyle is 28 years older than Sophia. in 9 years the sum of their ages will be 90. How old are they now?
Answer by ikleyn(52802)

.
K - S = 28.      (1)


The sum of their ages now is  90 - 2*9 = 72.  It gives you SECOND equation


K + S = 72.      (2)


Add equations (1) and (2). You will get


2K = 28 + 72 = 100  ===============>  K = 100/2 = 50.


Hence, S = 72 - 50 = 22.


ANSWER.  K is 50 years old;  S is 22 years old.